OpenStudy (anonymous):

Two players roll a 3-sided dice. The player who fails to get better rolls than previous ones loses. How likely is it for the first player to win?

OpenStudy (anonymous):

Is it 0.666666 ?

OpenStudy (anonymous):

Or 0.7037

OpenStudy (anonymous):

@satellite73 if u r finished there PLZ see this

OpenStudy (anonymous):

@lgbasallote can also do probability

OpenStudy (lgbasallote):

i highly doubt he can

OpenStudy (btaylor):

what does a 3-sided dice look like?

OpenStudy (anonymous):

i don't understand the question do they take turns?

OpenStudy (shane_b):

I believe they are taking turns...the first one to roll lower than the previous one's turn loses.

OpenStudy (anonymous):

I think if the first one roll is lower than the second one loses.

OpenStudy (anonymous):

What is the probability that the second roll will be higher than the 1st

OpenStudy (anonymous):

@BTaylor 3 sided dice looks like this.

OpenStudy (anonymous):

Just saying.

OpenStudy (btaylor):

if you list out the options: 1) Player 1 rolls a 1. Player 2 rolls either a 2 or a 3. If P2 rolls a 2, then P1 must roll a 3 to win. If P2 rolls a 3, then P1 loses. P=? 2) P1 rolls a 2. Player 2 must role a 3 to win. If P2 rolls a 3, then P1 loses. P=? 3) P1 rolls a 3. P2 automatically loses. (thanks @micahwood50 very nice)

OpenStudy (anonymous):

suppose the first player rolls a 2 and the second rolls a 2, what happens then ?

OpenStudy (anonymous):

First one wins

OpenStudy (anonymous):

"fails to get Higher"

OpenStudy (dumbcow):

so first player wins with these combinations: (3,1) (3,2) (2,1) that is probability of 3/9 or 1/3 chance player 1 wins after 1 turn

OpenStudy (anonymous):

ok so we can work in cases then right? we are only interested in the probability the first player wins

OpenStudy (anonymous):

(1,1) ,(2,1) ,(3,1) ,(2,2),(3,2) ,(3,3)

OpenStudy (anonymous):

rolls a 3 with probability \(\frac{1}{3}\) wins rolls a 2 with probability \(\frac{1}{3}\) player 2 rolls 1 or 2 with probability \(\frac{2}{3}\) wins rolls a 1 with probability \(\frac{1}{3}\) player 2 rolls a 1 with probability \(\frac{1}{3}\) wins

OpenStudy (anonymous):

Then it is 6/3^2 = 0.66666

OpenStudy (shane_b):

That makes sense...to me anyway.

OpenStudy (anonymous):

there are 18 total possibilities of Roll Number 2 in terms of Roll Number 1. Of those 18, 3 lead to number two winning. So the answer is 15/18

OpenStudy (anonymous):

unless i do not understand the question (which is possible) then it should be \[\frac{1}{33}+\frac{2}{9}+\frac{1}{9}\]

OpenStudy (anonymous):

oops typo

OpenStudy (anonymous):

U mean 1/3 + 2/9 + 1/9

OpenStudy (anonymous):

15/18 = 5/6

OpenStudy (anonymous):

\[\frac{1}{3}+\frac{2}{9}+\frac{1}{9}\]yeah but that is \(\frac{2}{3}\) which is not the answer given above

OpenStudy (anonymous):

2/3=0.666666

OpenStudy (anonymous):

so maybe i am wrong \[(3,1)(3,2)(3,3),(2,1)(2,2)(1,1) \] i still count 6 wins out of of 9

OpenStudy (anonymous):

For each of number one's rolls number 2 gets 3 rolls, that's 9 possibilities... of those 3 lead to winning so its 6/9 or 2/3. (i miscalculated earlier :/)

OpenStudy (anonymous):

So, the final answer is 2/3

OpenStudy (anonymous):

Ide bet my bippy its 2/3

OpenStudy (anonymous):

if i understand the question correctly, yes

OpenStudy (anonymous):

Thanx guys

OpenStudy (anonymous):

you bet your what? http://www.youtube.com/watch?v=HbvQa-7u3ss

OpenStudy (shane_b):

Just for kicks I did ran a simulation of 1000 games...player 1 won the game about 57-60% of the time (I ran the whole thing a few times).

OpenStudy (anonymous):

well theres your answer than 60%

OpenStudy (shane_b):

Being that the answer looked too low...I checked the code and found an error. Now I'm getting a solid 66% Player 1 win average across 1,000 test runs...with each test run averaging the results of 100 games. So yes, 66% must be correct :)